OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now?

OpenStudy (anonymous):

1) C = coach; T = Tim C = T + 10; C-3 = 2(T-3), so C = 2T - 6 + 3 = 2t - 3 2t - 3 = t + 10 t = 13, Tim's age now, so coach is 23; three years ago they were 10 & 20